USD Exchange Rate (USD/EUR)
% Cover Contracts Volume Options 1.01 1.22 1.48
0% 0% 25,000 100% 5,250 ((1.22-1.01)*25000) 0 -6,500 ((1.22-1.48)*2500)
0% 25% 25,000 75% 5,250 ((1.22-1.01)*25000) 0 -6,500 ((1.22-1.48)*2500)
0% 50% 25,000 50% 5,250 ((1.22-1.01)*25000) 0 -6,500 ((1.22-1.48)*2500)
0% 75% 25,000 25% 5,250 ((1.22-1.01)*25000) 0 -6,500 ((1.22-1.48)*2500)
0% 100% 25,000 0% 5,250 ((1.22-1.01)*25000) 0 -6,500 ((1.22-1.48)*2500)
25% 0% 25,000 100%
25% 25% 25,000 75%
25% 50% 25,000 50%
25% 75% 25,000 25%
25% 100% 25,000 0%
50% 0% 25,000 100%
50% 25% 25,000 75%
50% 50% 25,000 50%
50% 75% 25,000 25%
50% 100% 25,000 0%
75% 0% 25,000 100%
75% 25% 25,000 75%
75% 50% 25,000 50%
75% 75% 25,000 25%
75% 100% 25,000 0%
100% 0% 25,000 100% 3,725 (don't know how to get these numbers for 100% cover) -1,525 -1,525
100% 25% 25,000 75% 2,794 ( My professor gave them to me to go by but cant figure it out) -1,144 -1,144
100% 50% 25,000 50% 1,863 -763 -763
100% 75% 25,000 25% 931 -381 -381
100% 100% 25,000 0% 0

This Is for a hedging case I dont know how to get the rest of the numbers with different percent covers contracts and options please help. and put equations with them.

What would happen with a 100% hedge with forwards? A 100% hedge with options? Use the final sales volume of 25,000 and analyze the possible outcomes relative to the 'zero impact' scenario.
Answer #1

From the look and feel of the question and the numbers, we can safely assume that the company or the person in question has payable in EUR or it is a USD based company with volumes in EUR. So any upward movement is USDEUR is detrimental. Also, the cost of the options is 1,525. So, have tried to put equations to 100% scenario since you have specifically asked for the same. After we have put down equations, you can easily compare the scenarios with '0' hedge.

Also, if the numbers are assumed to be correct as specified in the question, here is the working. Kindly note that we have again assumed that your net gain / loss from futures contract is 0 due to the lack of any better information. Kindly provide additional details of the entire case.

USD Exchange Rate (USD/EUR)
% Cover Contracts Volume Options 1.01 1.22 1.48
100% 0% 25,000 100% =(25000*(1.22-1.01))-1525 =(25000*(1.22-1.22))-1525 Option shall not be exercised. Hence, outflow shall be limited to 1,525
100% 25% 25,000 75% =(25000*0.75*(1.22-1.01))-1144 You shall pay only 75% of 1,525 as a cost of the options. Hence, outflow shall be 1,144 You shall pay only 75% of 1,525 as a cost of the options. Hence, outflow shall be 1,144
100% 50% 25,000 50% =(25000*0.5*(1.22-1.01))-(0.5*1525) You shall pay only 50% of 1,525 as a cost of the options. Hence, outflow shall be 763 You shall pay only 50% of 1,525 as a cost of the options. Hence, outflow shall be 763
100% 75% 25,000 25% =(25000*0.25*(1.22-1.01))-(0.25*1525) You shall pay only 25% of 1,525 as a cost of the options. Hence, outflow shall be 381 You shall pay only 25% of 1,525 as a cost of the options. Hence, outflow shall be 381
100% 100% 25,000 0% No contracts available for the above sales volume No contracts available for the above sales volume No contracts available for the above sales volume